Question 1:

A) Suppose global climate models call for 2 °C of warming in the climate between 2020 and 2050, and that corn yields decrease 10% for every degree of warming. The baseline yield growth rate is 1% per year. What is the yield gain (in % terms) between 2020 and 2050 without climate change? What is the yield gain (in % terms) between 2020 and 2050 with climate change?

B) Now suppose that scientists have developed new, genetically modified corn varieties that can better tolerate heat and drought. These new varieties will be available for release in 2020. Suppose that these new varieties have the same yields as the old, traditional corn varieties, and the same baseline yield growth rates. Suppose that each degree of warming is only associated with 5% yield loss for the new varieties. What is the yield gain (in % terms) between 2020 and 2050 with these new varieties, assuming 2 °C of warming? How does this yield gain compare to yield gain under climate change with the old varieties?

C) Think about farmers’ decisions to substitute land between corn and wheat (for which no drought tolerant variety was released, but for which climate change is expected to decrease yields. What will happen to the area planted to wheat? Please justify your answer using concepts from class, and be very clear about the assumptions you use to arrive at your conclusion. You can depict it graphically if you wish.
Answer #1

A) yield gain (in % terms) between 2020 and 2050 without climate change=( 1% per year * 30) + (10% for every degree of warming *2) = 30 + 20 = 50% yield

Yield gain (in % terms) between 2020 and 2050 with climate change = ( 1% per year * 30) -(10% for every degree of warming *2) = 30 - 20 = 10% yield

B)yield gain (in % terms) between 2020 and 2050 with these new varieties, assuming 2 °C of warming = ( 1% per year * 30) -(5% for every degree of warming *2) = 30 - 10 = 20% yield

How does this yield gain compare to yield gain under climate change with the old varieties?

Answer: New yield = 20%; Old yield: 10%

C) By planting wheat, farmers will yield less than what they were yielding earlier by planting corns since drought tolerant variety of wheat was not invented.